OpenStudy (anonymous):

Solve the following system. 4x^2 + 9y^2 =72 x^2 - y^2 = 5 The solutions are: (), (), (), and () (remember to include the commas)

OpenStudy (anonymous):

@paki

OpenStudy (lolipop_drum312):

solving for y from eq1 when x = 3 4x^2 + 9y^2 = 72 4(3)^2 + 9y^2 = 72 36 + 9y^2 = 72 9y^2 = 72 - 36 9y^2 = 36 y^2 = 4 y = ± √(4) y = ± 2 y = 2, -2

OpenStudy (mrnood):

@Lolipop_Drum312 That does not look like a solution to me. You have chosen x=3 and x=-3 but not 'solved' for them

OpenStudy (lolipop_drum312):

Here is how you do it. Multiply the second equation by 9 to get 4x^2 + 9y^2 = 72 9x^2 - 9y^2 = 45 Add them to get 13x^2 = 117 ----> x^2 = 9 ----> x = +/- 3. Now substitute in the second original equation to get y

OpenStudy (mrnood):

That is correct

OpenStudy (lolipop_drum312):

ok good